# Bridging Gaps in Natural Language Processing for Yorùbá: A Systematic Review of a Decade of Progress and Prospects

### v1

@@ Abstract
Natural Language Processing (NLP) is becoming a dominant subset of artificial intelligence as the need to help machines understand human language looks indispensable. Several NLP applications are ubiquitous, partly due to the myriads of datasets being churned out daily through mediums like social networking sites. However, the growing development has not been evident in most African languages due to the persisting resource limitation, among other issues. Yorùbá language, a tonal and morphologically rich African language, suffers a similar fate, resulting in limited NLP usage. To encourage further research towards improving this situation, this systematic literature review aims to comprehensively analyse studies addressing NLP development for Yorùbá, identifying challenges, resources, techniques, and applications. A well-defined search string from a structured protocol was employed to search, select, and analyse 105 primary studies between $2014$ and $2024$ from reputable databases. The review highlights the scarcity of annotated corpora, limited availability of pre-trained language models, and linguistic challenges like tonal complexity and diacritic dependency as significant obstacles. It also revealed the prominent techniques, including rule-based methods, statistical methods, deep learning, and transfer learning, which were implemented alongside datasets of Yorùbá speech corpora, among others. The findings reveal a growing body of multilingual and monolingual resources, even though the field is constrained by socio-cultural factors such as code-switching and desertion of language for digital usage. This review synthesises existing research, providing a foundation for advancing NLP for Yorùbá and in African languages generally. It aims to guide future research by identifying gaps and opportunities, thereby contributing to the broader inclusion of Yorùbá and other under-resourced African languages in global NLP advancements.
@@ 1 Introduction
NLP techniques have witnessed fascinating evolution with diverse languages worldwide since its inception in the 1950s [19], initially with the rule-based approach. This involves defining linguistic rules for core language concepts, including semantics, pragmatics, morphology and phonology. These techniques have metamorphosed from the rule-based approach to statistical approaches, subsequently incorporating machine learning techniques, which utilise large-scale linguistic resources. Undoubtedly, the explosion of large datasets, mainly through social networking sites, necessitated the implementation of more advanced techniques. This births the introduction of deep learning methods [20], built on neural networks. Interestingly, the field of NLP witnessed revolutionisation through the advancement in deep learning, with notable models like the recurrent neural networks (RNNs), long short-term memory (LSTM), and more recently, the transformer architectures [21]. Moreover, this development has improved NLP performance in machine translation and question-answering tasks. More importantly, it gives way to a better understanding of sequential dependencies in natural language [22].
Furthermore, self-attention mechanisms and bidirectional training in models such as bidirectional encoder representations from transformers (BERT), alongside the generative pre-trained transformer (GPT-3) model developed by OpenAI [^1] https://openai.com/, resulted in remarkable improvement in natural language understanding [23]. This innovation continues to set the pace for the emergence of many improved large language models (LLMs) like the Large Language Model Architecture (LLaMA) [24], Mistral [25], and others. LLMs, usually trained on huge datasets, have been recorded to have achieved state-of-the-art result performance across various tasks, which transitions from task-specific to task-independent architectures [26]. However, the dataset need of these models translates into a boon and bane for NLP in the context of many under-resourced languages in the world. These languages, because they are less digitised, sparingly taught, harbour resource scarcity & low density, are less privileged, among other identifiers, are referred to as low-resource languages (LRLs) [27, 28].
@@
To promote NLP research involving Yorùbá language, it is essential to access the current status of such research involvement. To the best of our knowledge, no publication has addressed this need. Consequently, this research seeks to provide a systematic literature review (SLR) by synthesising existing research efforts in NLP for Yorùbá language. The study seeks to identify trends, highlight gaps, and propose directions for future research by answering specific research questions on the tasks addressed, methods employed, available resources, and prevailing challenges in NLP research involving Yorùbá language, either as a monolingual or bilingual component or as part of a multilingual study, where it is greatly emphasised. These questions are explicitly outlined in Section 3.2.
Ultimately, the motivation for this study is twofold. Firstly, it aspires to advocate for NLP solutions tailored to African languages, which remain underrepresented in global research. Ultimately, it is intended to bridge the gap between linguistic features unique to Yorùbá and their NLP representations, thereby contributing to linguistic preservation and technological inclusion as part of the Sustainable Development Goals (SDGs) of the United Nations (UN). The outcomes of this review are expected to serve as a comprehensive resource for researchers and practitioners working on Yorùbá language. Moreover, the findings would inform efforts to develop robust NLP systems accounting for the linguistic and cultural needs of Yorùbá while improving cross-lingual applications among other African languages.
@@ 2 Background of study
This section briefly introduces the Yorùbá language by discussing its constituents, such as letters and types, and detailing its root regarding its language family.
@@ 3.1 Review Planning
A systematic literature review requires formal planning and conscientiousness. Consequently, a defined protocol was followed to ensure the success of the entire process. The data, in the form of research articles, the review objectives, and defined selection criteria, were managed using the Covidence [^3] https://www.covidence.org/ platform, a software tool for effectively organising systematic reviews. It allows for a sequential flow, following a standard review protocol. Moreover, Zotero [^4] https://www.zotero.org/ is the reference manager, which ensures optimality and efficiency in the selection and writing process. Combining these tools with Microsoft Excel, the study benefits from tools that ensure flexibility and convenience, minimizing the required rigour of a typical systematic review.
@@
Following the review protocol shown in Figure 2, the first task was to define the research objectives and develop the research questions based on the objectives. Thereafter, the search strategy was developed. This includes defining the search string and the scope of the search. Subsequently, potential papers are searched through identified databases and screened based on the selection criteria. A quality assessment is performed on the included paper to ensure standard review and that the relevant data are extracted thereafter. Furthermore, the available data are analysed to obtain meaningful results reported comprehensively afterward. The next sections explained the processes in more detail.
@@ 3.2 Research Questions
Research questions were developed to define a precise template for the broad study objectives. Four questions have been carefully defined to cater to all possible domains intertwined with the research objectives by forming a structured set of interrogative statements whose answers are meant to provide insight into the research goal. The questions are motivated by the need to know the current status of Yorùbá language involvement in NLP by investigating the specific tasks, techniques, resources, and challenges. The research questions are presented sequentially in Table 1 to cater to structure flow and a quick overview.
@@
S/N Research Questions
RQ1 What NLP tasks have been addressed for Yorùbá language?
RQ2 What techniques have been employed for Yorùbá NLP?
RQ3 What language resources are available for Yorùbá language?
RQ4 What challenges are associated with NLP development involving Yorùbá language?
@@ 3.3 Search Strategy
This section describes the gathering of the vital primary studies and the “systematic” steps toward achieving the goal. It is a crucial phase planned to eliminate potential bias and incorporate randomisation in the studies’ selection and sample size determination. Reputable databases related to the subject matter were painstakingly explored to obtain all relevant studies for the systematic review using a well-defined search strategy. Generally, the central goal of the strategy involves attracting studies that have applied computational NLP approaches involving Yorùbá language, be it as a monolingual, bilingual, or multilingual NLP set-up.
@@ 3.3.1 Method and Scope of Search
The strategy initially employed an automated search method to obtain studies relevant to the systematic review. This method involved combing each electronic database with the defined keywords, boolean operators, and wild cards—as and when due. Furthermore, to ensure a representative sample and a higher recall, $9$ databases were targeted in total viz: Web of Science [^5] https://www.webofscience.com/, ScienceDirect [^6] https://www.sciencedirect.com/, Google Scholar [^7] https://scholar.google.com, Association for Computing Machinery (ACM) Digital Library [^8] https://dl.acm.org/, Institute of Electrical and Electronics Engineers (IEEE) Xplore [^9] https://ieeexplore.ieee.org/, Semantic Scholar [^10] https://www.semanticscholar.org/, Scopus [^11] https://www.scopus.com/, ScienceDirect [^12] https://www.sciencedirect.com/, and Association for Computational Linguistics (ACL) Anthology [^13] https://aclanthology.org/.
Moreover, with the motive of exploring a decade of the progress of NLP research in the Yorùbá language paradigm, studies between $2014$ and $2024$ were chosen. This decade range was also established by considering the time of writing this paper, and it is plausible since the studies of $2014$ were published towards the year’s end. In addition, to ensure potent quality, only peer-reviewed journal articles and conference papers were included in the search results, with the language of writing exclusively in English. Also, due to the computational requirement of NLP research, studies mainly in Computer Science, Engineering, and Computational Linguistics were considered, further focusing strictly on empirical studies. Figure 3 is a Pareto plot used to show the statistics of the initial search results across the databases, alongside the cumulative percentage. It shows that most papers were from Google Scholar, accounting for about $60\%$ of the search results, while the least was from IEEE Xplore. The curve shows the cumulative percentage of studies from each database. It indicates the percentage contribution of each database, with the least being IEEE Xplore.
@@ 3.3.2 Search Strings
The search strings denote the combination of keywords used in exploring the databases to extract potential primary studies. Potential keywords in the systematic review objectives were identified and combined to build this. However, before the combination, a few related research publications were scanned through to see the usage of the relevant phrases and words. Thus, since the study seeks to synthesize information vis-a-vis NLP research and solutions associated with Yorùbá language, the general search string is formulated based on the three segments and presented as follows:
@@
However, it is pertinent to mention that the search string was modified slightly for different databases based on their unique requirements and search functionalities.
@@ 3.3.3 Study Selection Criteria
Having obtained a substantial number of studies considered relevant to the study, the criteria to painstakingly select the most essential and directly aligned studies were defined. This involves a set of defined statements based on the research objectives and scope, making up the inclusion and exclusion criteria. It accounts for the desired year range, publication type, access to full-text documents, etc. Table 2 shows the inclusion and exclusion criteria, which are defined following the PICO criteria: population, intervention, context, and outcome [46]. In addition, Study Range and Publication Type and Publication Language were used to represent the publication years considered for the primary, the type of publication, and the language of publication, respectively. Furthermore, the two classes of criteria—inclusion and exclusion—defined statements to determine if the particular study is to be part of the review or irrelevant to it, respectively. Relevant articles were searched and gathered via the listed databases, bearing in mind the defined criteria.
Initially, $3144$ studies were obtained across the $9$ databases. The RIS files containing these studies from their respective sources were uploaded on the Covidence software tool. Due to the inevitable intersection across databases, the software detected a total duplicate count of $816$, while $50$ duplicates were manually detected. Again, following the selection criteria, abstract screening was initially carried out, leaving out a total of $1865$ from the unique $2278$ studies. Moreover, full-text screening was carried out on the remaining $413$ studies, and a total of $308$ studies were discarded based on defined reasons. Eventually, $105$ primary studies were obtained for inclusion in the review.
Subsequently, backward and forward snowballing techniques were also used to capture studies that might have been missed in database searches due to conflicting terminology, following [48]. For backward snowballing, reference lists of a quarter of the $105$ articles were surveyed to check for additional relevant studies. It was observed that the identified publications were part of the initial search results across databases. Additionally, forward snowballing, which involves checking for publications that have cited the potential primary study, was carried out. Here, $5$ pre-prints and $1$ review publications were discovered; however, they were not included due to the defined exclusion criteria. The snowballing loops were ended after no new studies were obtained. Ultimately, $105$ primary studies were included in the SLR after conducting a quality assessment check, presented in Section 3.4. The whole selection process ensures a detailed and comprehensive overview by incorporating the Preferred Reporting Items for Systematic Reviews and Meta-Analysis (PRISMA) framework, which is also presented in Figure 4, following [47].
@@
Study Range Publications between $2014$ $2024$ inclusive Publications before $2014$
@@ 3.4 Quality Assessment
This section shows further efforts to ensure the credibility of primary studies used in the systematic literature review. Specific quality questions were used for verification to further ascertain the reliability of the primary studies to be eventually included in the SLR. Thus, a quality checklist was created, following the guidelines in [46]. The response associated with each question is defined through a multiple-choice response: Yes, Partially, and No. A “Yes” response shows higher certainty, thus getting a score of $2$. Similarly, a “Partially” response is assigned $1$, while a “No” response is assigned 00. For this research purpose, a total quality score below the median score is to be excluded. The median is calculated as the value in the $\Big{(}\frac{n+1}{2}\Big{)}^{\text{th}}$ position, where $n$ is the maximum quality score, which equals $16$. Thus, for quality scores $S_{i}=\{1,2,\dots,16\}$, the median position, $M_{p}$, is derived as follows:
@@
Consequently, a paper with a score less than $8.5$ out of the $16$ total score will be excluded. However, no study falls in this range. Hence, the total $105$ primary studies are included in the systematic review. Table 3 shows the checklist containing the quality domain and the questions asked for each.
@@
Methodology and study design Are the study’s methods and experimental design clearly defined?
@@
After assessing each potential primary study based on the quality questions, scores were assigned—a maximum of 16 for each. The results are plotted through the histogram in Figure 5. The plot shows that most of the primary studies have a score of $15$, followed by $14$. Moreover, only a minute quantity of the primary studies has a total score between $9$ and $12$, with $9$ being the minimum total score being $9$, while the maximum is $16$. This shows high quality in the included primary studies, potentially making this review’s results highly valuable.

### v2

@@ Abstract
Natural Language Processing (NLP) is becoming a dominant subset of artificial intelligence, as the need for machines to understand human language becomes increasingly indispensable. Several NLP applications are ubiquitous, partly due to the myriad datasets being churned out daily through mediums like social networking sites. However, the growing development has not been evident in most African languages due to persistent resource limitations, among other issues. Yorùbá language, a tonal and morphologically rich African language, suffers a similar fate, resulting in limited NLP usage. To encourage further research towards improving this situation, this systematic literature review aims to comprehensively analyse studies addressing NLP development for Yorùbá, identifying challenges, resources, techniques, and applications. A well-defined search string from a structured protocol was employed to search, select, and analyse 105 primary studies between $2014$ and $2024$ from reputable databases. The review highlights the scarcity of annotated corpora, the limited availability of pre-trained language models, and linguistic challenges like tonal complexity and diacritic dependency as significant obstacles. It also revealed the prominent techniques, including rule-based methods, statistical methods, deep learning, and transfer learning, which were implemented alongside datasets of Yorùbá speech corpora, among others. The findings reveal a growing body of multilingual and monolingual resources, even though the field is constrained by socio-cultural factors such as code-switching and the desertion of language for digital usage. This review synthesises existing research, providing a foundation for advancing NLP for Yorùbá and in African languages generally. It aims to guide future research by identifying gaps and opportunities, thereby contributing to the broader inclusion of Yorùbá and other under-resourced African languages in global NLP advancements.
@@ 1 Introduction
NLP techniques have witnessed fascinating evolution with diverse languages worldwide since their inception in the 1950s [74], initially with the rule-based approach. This involves defining linguistic rules for core language concepts, including semantics, pragmatics, morphology and phonology. These techniques have metamorphosed from the rule-based approach to statistical approaches, subsequently incorporating machine learning techniques, which utilise large-scale linguistic resources. Undoubtedly, the explosion of large datasets, mainly through social networking sites, necessitated the implementation of more advanced techniques. This gives rise to the introduction of deep learning methods [79], built on neural networks. Interestingly, the field of NLP witnessed revolutionisation through the advancement in deep learning, with notable models like the recurrent neural networks (RNNs), long short-term memory (LSTM), and more recently, the transformer architectures [76]. Moreover, this development has improved NLP performance in machine translation and question-answering tasks. More importantly, it gives way to a better understanding of sequential dependencies in natural language [120].
Furthermore, self-attention mechanisms and bidirectional training in models such as bidirectional encoder representations from transformers (BERT), alongside the generative pre-trained transformer (GPT-3) model developed by OpenAI [^1] https://openai.com/, resulted in remarkable improvement in natural language understanding [78]. This innovation continues to set the pace for the emergence of many improved large language models (LLMs) like the Large Language Model Architecture (LLaMA) [119], Mistral [73], and others. LLMs, usually trained on huge datasets, have been recorded to have achieved state-of-the-art result performance across various tasks, which transitions from task-specific to task-independent architectures [103]. However, the huge dataset needs of these models translate into a boon and bane for NLP in the context of many under-resourced languages in the world. These languages, because they are less digitised, sparingly taught, harbour resource scarcity & low density, are less privileged, among other identifiers, are referred to as low-resource languages (LRLs) [114, 6].
@@
To promote NLP research involving Yorùbá language, it is essential to examine the current landscape of research efforts, identify existing gaps, and evaluate available tools and resources. While some work has been done in isolated tasks such as diacritic restoration, machine translation, and sentiment analysis, there is no comprehensive review that brings together these fragmented efforts. As a result, researchers and developers lack a centralised knowledge base to inform further progress in Yorùbá NLP. This lack of synthesis impedes collaboration, resource-sharing, and efficient advancement in the field. This study, therefore, aims to fill this critical gap by conducting a systematic literature review (SLR) that synthesises existing research in NLP for the Yorùbá language. Specifically, the review seeks to:
  • Identify the range of NLP tasks addressed in the literature,
  • Categorise the methods and techniques employed,
  • Document the existing resources and tools, and
  • Highlight the unique linguistic challenges and how they have been addressed.
These insights will offer a clear map of what has been achieved so far, where the limitations lie, and where future research should be directed. The guiding research questions for this review are explicitly outlined in Section 3.2.
The intended contribution of this review is twofold. Firstly, it serves as a foundational reference for researchers, developers, and language technologists interested in Yorùbá and other low-resource African languages, by offering a structured overview of existing work and identifying key gaps and opportunities. Secondly, by revealing progress and bottlenecks, this review can support targeted development of language tools, datasets, and models tailored to the linguistic features of Yorùbá. In doing so, it contributes to the broader goals of linguistic preservation, technological inclusion, and equitable representation of African languages in NLP, which is an effort supporting the United Nations’ Sustainable Development Goals (SDGs). Ultimately, the outcomes of this review are expected to guide future research agendas, inform policy and funding decisions, and foster collaborations toward inclusive NLP development.
@@ 2 Background of study
This section briefly introduces the Yorùbá language by discussing its constituents, such as letters and types, and detailing its roots regarding its language family.
@@ 3.1 Review Planning
A systematic literature review requires formal planning and meticulous attention to detail. Consequently, a defined protocol was followed to ensure the success of the entire process. To minimise the manual rigour typically required in systematic reviews, three key tools were employed: Covidence [^3] https://www.covidence.org/, Zotero [^4] https://www.zotero.org/, and Microsoft Excel. Covidence streamlined multiple stages of the review process. It enabled the bulk import of studies in RIS format and automatically detected and removed duplicate records from various article databases, eliminating the need for manual cross-checking. Covidence also facilitated the title/abstract and full-text screening phases through a structured interface that supported inclusion/exclusion tagging and decision logging by multiple reviewers. Zotero served as the reference management platform, supporting efficient bibliographic organisation, duplicate checking, and citation formatting. Microsoft Excel was used to define a structured data extraction template and to streamline the quality assessment phase, thereby reducing the likelihood of manual errors.
These tools collectively reduced human effort in screening, deduplication, tagging, and data organisation. Furthermore, they addressed methodological bottlenecks, including inconsistent inclusion decisions, manual duplicate removal, error-prone data tracking, and fragmented reviewer collaboration. Ultimately, they helped in improving reproducibility, transparency, and efficiency, as well as the adoption of a more reliable and scalable review workflow.
@@
Moreover, following the review protocol shown in Figure 2, the first task was to define the research objectives and develop the research questions based on the objectives. Thereafter, the search strategy was developed. This includes determining the search string and the scope of the search. Subsequently, potential papers are searched through identified databases and screened based on the selection criteria. A quality assessment is performed on the included paper to ensure standard review and that the relevant data are extracted thereafter. Furthermore, the available data are analysed to obtain the results reported comprehensively afterward. The following sections provide a more detailed explanation of the processes.
@@ 3.2 Research Questions
Research questions were developed to define a precise template for the broad study objectives. Four research questions presented in Table 1 were carefully formulated to guide a structured and focused review of the current state of NLP research involving the Yorùbá language. Each question targets a distinct dimension of the NLP pipeline and contributes to achieving the study’s overall objective: to investigate the development, resources, methods, and challenges in Yorùbá NLP. RQ1 helps to identify the variety of computational tasks that have received attention in Yorùbá NLP research. It is meant to establish the breadth of work done and uncover underexplored tasks. RQ2 focuses on the computational methods and frameworks used across these studies. This helps to track methodological trends and possibly evaluate how well existing approaches align with the linguistic complexities of Yorùbá. Additionally, RQ3 aims to assess the existing resources and ultimately identify resource gaps that hinder progress. Finally, RQ4 addresses the core limitations and bottlenecks that researchers face, helping to understand the challenges that are vital for proposing actionable solutions and advancing future research in various domains.
Collectively, these research questions seek to provide a well-encompassing basis for analysing the current state of Yorùbá NLP, highlighting strengths, gaps, and priorities for further investigation, directly supporting the main objectives of the study.
@@
$RQ_{i}$ Research Questions
$RQ_{1}$ What NLP tasks have been addressed for Yorùbá language?
$RQ_{2}$ What techniques have been employed for Yorùbá NLP?
$RQ_{3}$ What language resources are available for Yorùbá language?
$RQ_{4}$ What challenges are associated with NLP development involving Yorùbá language?
@@ 3.3 Search Strategy
This section describes the gathering of the vital primary studies and the systematic steps toward achieving the goal. This is a crucial phase designed to eliminate potential bias and incorporate randomization in the selection and sample size determination. Reputable databases related to the subject matter were thoroughly examined to identify all relevant studies for the systematic review, utilizing a well-defined search strategy. Generally, the central goal of the strategy involves attracting studies that have applied computational NLP approaches involving Yorùbá language, be it as a monolingual, bilingual, or multilingual NLP set-up.
@@ 3.3.1 Method and Scope of Search
The strategy initially employed an automated search method to obtain studies relevant to the systematic review. This method involved digging into each electronic database with the defined keywords, boolean operators, and wild cards—as and when due. Furthermore, to ensure a representative sample and a higher recall, $9$ databases were targeted in total, viz: Web of Science [^5] https://www.webofscience.com/, ScienceDirect [^6] https://www.sciencedirect.com/, Google Scholar [^7] https://scholar.google.com, Association for Computing Machinery (ACM) Digital Library [^8] https://dl.acm.org/, Institute of Electrical and Electronics Engineers (IEEE) Xplore [^9] https://ieeexplore.ieee.org/, Semantic Scholar [^10] https://www.semanticscholar.org/, Scopus [^11] https://www.scopus.com/, ScienceDirect [^12] https://www.sciencedirect.com/, and Association for Computational Linguistics (ACL) Anthology [^13] https://aclanthology.org/.
Moreover, with the motive of exploring a decade of the progress of NLP research in the Yorùbá language paradigm, studies between $2014$ and $2024$ were chosen. This decade range was also established by considering the time of writing this paper, and it is plausible since the studies of $2014$ were published towards the year’s end. In addition, to ensure potent quality, only peer-reviewed journal articles and conference papers were included in the search results, with the language of writing exclusively in English. Additionally, due to the computational requirements of NLP research, studies primarily in Computer Science, Engineering, and Computational Linguistics were considered, with a strict focus on empirical studies. Figure 3 is a Pareto plot used to show the statistics of the initial search results across the databases, alongside the cumulative percentage. It shows that most papers were from Google Scholar, accounting for about $60\%$ of the search results, while the least were from IEEE Xplore. The curve shows the cumulative percentage of studies from each database. It indicates the percentage contribution of each database, with the least being IEEE Xplore.
@@ 3.3.2 Search Strings
The search strings denote the combination of keywords used in exploring the databases to extract potential primary studies. Potential keywords in the systematic review objectives were identified and combined to build this. However, before the combination, a few related research publications were scanned through to see the usage of the relevant phrases and words. Thus, since the study seeks to synthesize information vis-à-vis NLP research and solutions associated with Yorùbá language, the general search string is formulated based on the three segments and presented as follows:
@@
However, it is pertinent to mention that the search string was modified slightly for different databases based on their unique requirements and search functionalities, and the domains were restricted to computer science and computational linguistics. Moreover, to obtain search records from Google Scholar, a software tool named “Publish or Perish”, which obtains and analyzes scholarly citations, was used. This tool leverages many data sources, including Google Scholar, to get the raw citations seamlessly.
@@ 3.3.3 Study Selection Criteria
Having obtained a substantial number of studies considered relevant to the study, the criteria to painstakingly select the most essential and directly aligned studies were defined. This involves a set of defined statements based on the research objectives and scope, making up the inclusion and exclusion criteria. It accounts for the desired year range, publication type, access to full-text documents, etc. Table 2 shows the inclusion and exclusion criteria, which are defined following the PICO criteria: population, intervention, context, and outcome [77]. In addition, Study Range and Publication Type, and Publication Language were used to represent the publication years considered for the primary, the type of publication, and the language of publication, respectively. Furthermore, the two classes of criteria—inclusion and exclusion—defined statements to determine if the particular study is to be part of the review or irrelevant to it, respectively. Relevant articles were searched and gathered via the listed databases, bearing in mind the defined criteria.
@@
Study Range Publications between $2014$ & $2024$ inclusive Publications before $2014$
@@
Initially, $3144$ studies were obtained across the $9$ databases. The RIS files containing these studies from their respective sources were uploaded to the Covidence software tool. Due to the inevitable intersection across databases, the software detected a total duplicate count of $816$, while $50$ duplicates were manually detected. Again, following the selection criteria, abstract screening was initially carried out, leaving out a total of $1865$ from the unique $2278$ studies. Moreover, full-text screening was carried out on the remaining $413$ studies, and a total of $308$ studies were discarded based on defined reasons. Eventually, $105$ primary studies were obtained for inclusion in the review.
Subsequently, backwards and forward snowballing techniques were also used to capture studies that might have been missed in database searches due to conflicting terminology, following [125]. For backward snowballing, reference lists of a quarter of the $105$ articles were surveyed to check for additional relevant studies. It was observed that the identified publications were part of the initial search results across databases. Additionally, forward snowballing, which involves checking for publications that have cited the potential primary study, was carried out. Here, $5$ pre-prints and $1$ review publications were discovered; however, they were not included due to the defined exclusion criteria. The snowballing loops were ended after no new studies were obtained. Ultimately, $105$ primary studies were included in the SLR after conducting a quality assessment check, presented in Section 3.4. The whole selection process ensures a detailed and comprehensive overview by incorporating the Preferred Reporting Items for Systematic Reviews and Meta-Analysis (PRISMA) framework, which is also presented in Figure 4, following [102].
@@ 3.4 Quality Assessment
This section shows further efforts to ensure the credibility of primary studies used in the systematic literature review. Specific quality questions were used for verification to further ascertain the reliability of the primary studies to be eventually included in the SLR. Thus, a quality checklist was created, following the guidelines in [77]. The response associated with each question is defined through a multiple-choice response: Yes, Partially, and No. A “Yes” response shows higher certainty, thus getting a score of $2$ . Similarly, a “Partially” response is assigned $1$, while a “No” response is assigned 0. For this research purpose, a total quality score below the median score is to be excluded. The median is calculated as the value in the $\Big(\frac{n+1}{2}\Big)^{\text{th}}$ position, where $n$ is the maximum quality score, which equals $16$. Thus, for quality scores $S_{i}=\{1,2,\dots,16\}$, the median position, $M_{p}$, is derived as follows:
@@
Consequently, a paper with a score less than $8.5$ out of the $16$ total score will be excluded. However, no study falls in this range. Hence, the total $105$ primary studies are included in the systematic review. The quality assessment technique is based on the study by [122] and was extended using a measure of central tendency, the median, as it accurately measures the center of a dataset’s distribution and is also not affected by outliers. Table 3 shows the checklist containing the quality domain and the questions asked for each.
@@
Methodology & study design Are the study’s methods and experimental design clearly defined?
@@
After assessing each potential primary study based on the quality questions, scores were assigned—a maximum of 16 for each. The results are plotted through the histogram in Figure 5. The plot shows that most of the primary studies have a score of $15$, followed by $14$. Moreover, only a minute quantity of the primary studies have a total score between $9$ and $12$, with $9$ being the minimum total score being $9$, while the maximum is $16$. This shows high quality in the included primary studies, potentially making this review’s results highly valuable.